OpenStudy (anonymous):

What type of triangle is formed by joining the points D(7, 3), E(8, 1), and F(4, -1)?

OpenStudy (anonymous):

help? @dan815

OpenStudy (er.mohd.amir):

find distance of sides using d*d=(x2-x1)^2+(y2-y1)^2

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

these are my choices equilateral triangle isosceles triangle right triangle acute scalene triangle obtuse scalene triangle

OpenStudy (er.mohd.amir):

if all sides equal then equilateral if two sides equal isosceles a^2=b^2+c^2 then right angle